\begin{thebibliography}{10}

\bibitem{Ahmadi:2016:NFE:2857705.2857713}
Mansour Ahmadi, Dmitry Ulyanov, Stanislav Semenov, Mikhail Trofimov, and
  Giorgio Giacinto.
\newblock Novel feature extraction, selection and fusion for effective malware
  family classification.
\newblock In {\em Proceedings of the Sixth ACM Conference on Data and
  Application Security and Privacy}, CODASPY '16, pages 183--194, New York, NY,
  USA, 2016. ACM.
\newblock URL: \url{http://doi.acm.org/10.1145/2857705.2857713}, \href
  {http://dx.doi.org/10.1145/2857705.2857713}
  {\path{doi:10.1145/2857705.2857713}}.

\bibitem{10.1007/978-3-319-51966-1_17}
Saed Alrabaee, Paria Shirani, Mourad Debbabi, and Lingyu Wang.
\newblock On the feasibility of malware authorship attribution.
\newblock In Fr{\'e}d{\'e}ric Cuppens, Lingyu Wang, Nora Cuppens-Boulahia,
  Nadia Tawbi, and Joaquin Garcia-Alfaro, editors, {\em Foundations and
  Practice of Security}, pages 256--272, Cham, 2016. Springer International
  Publishing.

\bibitem{Bagga2017}
Naman Bagga.
\newblock Measuring the effectiveness of generic malware models, 2017.

\bibitem{10.1007/978-3-319-60876-1_10}
Thomas Barabosch, Niklas Bergmann, Adrian Dombeck, and Elmar Padilla.
\newblock Quincy: Detecting host-based code injection attacks in memory dumps.
\newblock In Michalis Polychronakis and Michael Meier, editors, {\em Detection
  of Intrusions and Malware, and Vulnerability Assessment}, pages 209--229,
  Cham, 2017. Springer International Publishing.

\bibitem{DBLP:journals/corr/BhattacharyaMBC16}
Sukriti Bhattacharya, H{\'{e}}ctor~D. Men{\'{e}}ndez, Earl~T. Barr, and David
  Clark.
\newblock Itect: Scalable information theoretic similarity for malware
  detection.
\newblock {\em CoRR}, abs/1609.02404, 2016.
\newblock URL: \url{http://arxiv.org/abs/1609.02404}, \href
  {http://arxiv.org/abs/1609.02404} {\path{arXiv:1609.02404}}.

\bibitem{french2015}
Philippe Biondi, Xavier Mehrenberger, and Sarah Zennou.
\newblock Rebus : un bus de communication facilitant la coopération entre
  outils d’analyse de sécurité.
\newblock In {\em Symposium on Information and Communications Security}, 2015.

\bibitem{Borbely2016}
Rebecca~Schuller Borbely.
\newblock On normalized compression distance and large malware.
\newblock {\em Journal of Computer Virology and Hacking Techniques},
  12(4):235--242, Nov 2016.
\newblock URL: \url{https://doi.org/10.1007/s11416-015-0260-0}, \href
  {http://dx.doi.org/10.1007/s11416-015-0260-0}
  {\path{doi:10.1007/s11416-015-0260-0}}.

\bibitem{7836677}
E.~Burnaev and D.~Smolyakov.
\newblock One-class svm with privileged information and its application to
  malware detection.
\newblock In {\em 2016 IEEE 16th International Conference on Data Mining
  Workshops (ICDMW)}, pages 273--280, Dec 2016.
\newblock \href {http://dx.doi.org/10.1109/ICDMW.2016.0046}
  {\path{doi:10.1109/ICDMW.2016.0046}}.

\bibitem{russian2016}
Evgeny Burnayev and Dmitry Smolyakov.
\newblock One-class machine of reference vectors using privileged information.
\newblock In {\em Information technologies and systems (ITaS)}, 2016.

\bibitem{Celik:2017:FCP:3041008.3041018}
Z.~Berkay Celik, Patrick McDaniel, and Rauf Izmailov.
\newblock Feature cultivation in privileged information-augmented detection.
\newblock In {\em Proceedings of the 3rd ACM on International Workshop on
  Security And Privacy Analytics}, IWSPA '17, pages 73--80, New York, NY, USA,
  2017. ACM.
\newblock URL: \url{http://doi.acm.org/10.1145/3041008.3041018}, \href
  {http://dx.doi.org/10.1145/3041008.3041018}
  {\path{doi:10.1145/3041008.3041018}}.

\bibitem{DBLP:journals/corr/CelikMIPS16}
Z.~Berkay Celik, Patrick~D. McDaniel, Rauf Izmailov, Nicolas Papernot, and
  Ananthram Swami.
\newblock Building better detection with privileged information.
\newblock {\em CoRR}, abs/1603.09638, 2016.
\newblock URL: \url{http://arxiv.org/abs/1603.09638}, \href
  {http://arxiv.org/abs/1603.09638} {\path{arXiv:1603.09638}}.

\bibitem{7723750}
A.~Dinh, D.~Brill, Y.~Li, and W.~He.
\newblock Malware sequence alignment.
\newblock In {\em 2016 IEEE International Conferences on Big Data and Cloud
  Computing (BDCloud), Social Computing and Networking (SocialCom), Sustainable
  Computing and Communications (SustainCom) (BDCloud-SocialCom-SustainCom)},
  pages 613--617, Oct 2016.
\newblock \href
  {http://dx.doi.org/10.1109/BDCloud-SocialCom-SustainCom.2016.96}
  {\path{doi:10.1109/BDCloud-SocialCom-SustainCom.2016.96}}.

\bibitem{7527757}
J.~Drew, T.~Moore, and M.~Hahsler.
\newblock Polymorphic malware detection using sequence classification methods.
\newblock In {\em 2016 IEEE Security and Privacy Workshops (SPW)}, pages
  81--87, May 2016.
\newblock \href {http://dx.doi.org/10.1109/SPW.2016.30}
  {\path{doi:10.1109/SPW.2016.30}}.

\bibitem{Drew2017}
Jake Drew, Michael Hahsler, and Tyler Moore.
\newblock Polymorphic malware detection using sequence classification methods
  and ensembles.
\newblock {\em EURASIP Journal on Information Security}, 2017(1):2, Jan 2017.
\newblock URL: \url{https://doi.org/10.1186/s13635-017-0055-6}, \href
  {http://dx.doi.org/10.1186/s13635-017-0055-6}
  {\path{doi:10.1186/s13635-017-0055-6}}.

\bibitem{8064195}
J.~Fan, C.~Guan, K.~Ren, Y.~Cui, and C.~Qiao.
\newblock Spabox: Safeguarding privacy during deep packet inspection at a
  middlebox.
\newblock {\em IEEE/ACM Transactions on Networking}, 25(6):3753--3766, Dec
  2017.
\newblock \href {http://dx.doi.org/10.1109/TNET.2017.2753044}
  {\path{doi:10.1109/TNET.2017.2753044}}.

\bibitem{Fields2016}
Tonya Fields and Jonathan Graham.
\newblock Classifying network attack data using random forest.
\newblock In {\em CATA}, Dec 2016.

\bibitem{Fowler2015}
Charles~A. Fowler.
\newblock {\em A HYBRID INTELLIGENCE/MULTI-AGENT SYSTEM FOR MINING INFORMATION
  ASSURANCE DATA}.
\newblock PhD thesis, May 2015.

\bibitem{DBLP:journals/corr/GaoWQ17}
Ji~Gao, Beilun Wang, Zeming Lin, and Weilin Xu~Yanjun Qi.
\newblock Deepcloak: Masking deep neural network models for robustness against
  adversarial samples.
\newblock {\em CoRR}, abs/1702.06763, 2017.
\newblock URL: \url{http://arxiv.org/abs/1702.06763}, \href
  {http://arxiv.org/abs/1702.06763} {\path{arXiv:1702.06763}}.

\bibitem{DBLP:journals/corr/GarciaM16}
Felan Carlo~C. Garcia and Felix P.~Muga II.
\newblock Random forest for malware classification.
\newblock {\em CoRR}, abs/1609.07770, 2016.
\newblock URL: \url{http://arxiv.org/abs/1609.07770}, \href
  {http://arxiv.org/abs/1609.07770} {\path{arXiv:1609.07770}}.

\bibitem{10.1007/978-3-319-71246-8_3}
Severin Gsponer, Barry Smyth, and Georgiana Ifrim.
\newblock Efficient sequence regression by learning linear models in
  all-subsequence space.
\newblock In Michelangelo Ceci, Jaakko Hollm{\'e}n, Ljup{\v{c}}o Todorovski,
  Celine Vens, and Sa{\v{s}}o D{\v{z}}eroski, editors, {\em Machine Learning
  and Knowledge Discovery in Databases}, pages 37--52, Cham, 2017. Springer
  International Publishing.

\bibitem{Hansen2017}
Joachim Hansen.
\newblock The study of keyword search in open source search engines and digital
  forensics tools with respect to the needs of cyber crime investigations,
  2017.

\bibitem{Hassen2017}
M.~Hassen, M.~Carvalho, and P.~Chan.
\newblock Malware classification using static analysis based features.
\newblock In {\em IEEE Symposium on Computational Intelligence in Cyber
  Security}, Dec 2017.

\bibitem{2018arXiv180204365H}
M.~{Hassen} and P.~K. {Chan}.
\newblock {Learning a Neural-network-based Representation for Open Set
  Recognition}.
\newblock {\em ArXiv e-prints}, February 2018.
\newblock \href {http://arxiv.org/abs/1802.04365} {\path{arXiv:1802.04365}}.

\bibitem{Hassen:2017:SFC:3029806.3029824}
Mehadi Hassen and Philip~K. Chan.
\newblock Scalable function call graph-based malware classification.
\newblock In {\em Proceedings of the Seventh ACM on Conference on Data and
  Application Security and Privacy}, CODASPY '17, pages 239--248, New York, NY,
  USA, 2017. ACM.
\newblock URL: \url{http://doi.acm.org/10.1145/3029806.3029824}, \href
  {http://dx.doi.org/10.1145/3029806.3029824}
  {\path{doi:10.1145/3029806.3029824}}.

\bibitem{10.1007/978-981-10-7080-8_17}
Tingting He, Jingfeng Xue, Jianwen Fu, Yong Wang, and Chun Shan.
\newblock Research on malicious code analysis method based on semi-supervised
  learning.
\newblock In Ming Xu, Zheng Qin, Fei Yan, and Shaojing Fu, editors, {\em
  Trusted Computing and Information Security}, pages 227--241, Singapore, 2017.
  Springer Singapore.

\bibitem{7523365}
X.~Hu, J.~Jang, T.~Wang, Z.~Ashraf, M.~P. Stoecklin, and D.~Kirat.
\newblock Scalable malware classification with multifaceted content features
  and threat intelligence.
\newblock {\em IBM Journal of Research and Development}, 60(4):6:1--6:11, July
  2016.
\newblock \href {http://dx.doi.org/10.1147/JRD.2016.2559378}
  {\path{doi:10.1147/JRD.2016.2559378}}.

\bibitem{fselection2018}
Seong~Oun Hwanga, Trong~Kha Nguyenb, and Vu~Duc Ly.
\newblock Feature selection for malware classification.
\newblock Technical report, January 2018.

\bibitem{imran2016}
Mohammad Imran.
\newblock {\em Evlauation of Hidden Markov Model for Malware Behavioural
  Classification}.
\newblock PhD thesis, 2016.

\bibitem{203684}
Roberto Jordaney, Kumar Sharad, Santanu~K. Dash, Zhi Wang, Davide Papini, Ilia
  Nouretdinov, and Lorenzo Cavallaro.
\newblock Transcend: Detecting concept drift in malware classification models.
\newblock In {\em 26th {USENIX} Security Symposium ({USENIX} Security 17)},
  pages 625--642, Vancouver, BC, 2017. {USENIX} Association.
\newblock URL:
  \url{https://www.usenix.org/conference/usenixsecurity17/technical-sessions/presentation/jordaney}.

\bibitem{8268747}
T.~M. Kebede, O.~Djaneye-Boundjou, B.~N. Narayanan, A.~Ralescu, and D.~Kapp.
\newblock Classification of malware programs using autoencoders based deep
  learning architecture and its application to the microsoft malware
  classification challenge (big 2015) dataset.
\newblock In {\em 2017 IEEE National Aerospace and Electronics Conference
  (NAECON)}, pages 70--75, June 2017.
\newblock \href {http://dx.doi.org/10.1109/NAECON.2017.8268747}
  {\path{doi:10.1109/NAECON.2017.8268747}}.

\bibitem{10.1007/978-981-10-7605-3_215}
Hae-Jung Kim.
\newblock Image-based malware classification using convolutional neural
  network.
\newblock In James~J. Park, Vincenzo Loia, Gangman Yi, and Yunsick Sung,
  editors, {\em Advances in Computer Science and Ubiquitous Computing}, pages
  1352--1357, Singapore, 2018. Springer Singapore.

\bibitem{Kim2017}
Jin-Young Kim, Seok-Jun Bu, and Sung-Bae Cho.
\newblock Malware detection using deep transferred generative adversarial
  networks.
\newblock In Derong Liu, Shengli Xie, Yuanqing Li, Dongbin Zhao, and
  El-Sayed~M. El-Alfy, editors, {\em Neural Information Processing}, pages
  556--564, Cham, 2017. Springer International Publishing.

\bibitem{korean2016}
Nak-Hyun Kim, Byung ik~Kim, and Tae jin Lee.
\newblock Performance analysis of the malware classification method in
  accordance with the changes in assembly code.
\newblock 2016.

\bibitem{2018arXiv180204528K}
F.~{Kreuk}, A.~{Barak}, S.~{Aviv-Reuven}, M.~{Baruch}, B.~{Pinkas}, and
  J.~{Keshet}.
\newblock {Adversarial Examples on Discrete Sequences for Beating Whole-Binary
  Malware Detection}.
\newblock {\em ArXiv e-prints}, February 2018.
\newblock \href {http://arxiv.org/abs/1802.04528} {\path{arXiv:1802.04528}}.

\bibitem{NODE07303202}
Hyun-Jong Lee, Heo-Jae Hyeok, and Doosung Hwang.
\newblock {Performance Comparison of Machine Learning Algorithms for Malware
  Detection}.
\newblock volume~{26}, pages 143--146. The Korean Society Of Computer And
  Information, 2018.
\newblock URL: \url{http://www.dbpia.co.kr/Article/NODE07303202}.

\bibitem{msrt18}
Microsoft.
\newblock Sam cybersecurity engagement kit, 2018.
\newblock URL:
  \url{https://assets.microsoft.com/en-nz/cybersecurity-sam-engagement-kit.pdf}.

\bibitem{7856826}
B.~N. Narayanan, O.~Djaneye-Boundjou, and T.~M. Kebede.
\newblock Performance analysis of machine learning and pattern recognition
  algorithms for malware classification.
\newblock In {\em 2016 IEEE National Aerospace and Electronics Conference
  (NAECON) and Ohio Innovation Summit (OIS)}, pages 338--342, July 2016.

\bibitem{8062202}
A.~P. Norton and Y.~Qi.
\newblock Adversarial-playground: A visualization suite showing how adversarial
  examples fool deep learning.
\newblock In {\em 2017 IEEE Symposium on Visualization for Cyber Security
  (VizSec)}, pages 1--4, Oct 2017.
\newblock \href {http://dx.doi.org/10.1109/VIZSEC.2017.8062202}
  {\path{doi:10.1109/VIZSEC.2017.8062202}}.

\bibitem{8102915}
R.~Paranthaman and B.~Thuraisingham.
\newblock Malware collection and analysis.
\newblock In {\em 2017 IEEE International Conference on Information Reuse and
  Integration (IRI)}, pages 26--31, Aug 2017.
\newblock \href {http://dx.doi.org/10.1109/IRI.2017.92}
  {\path{doi:10.1109/IRI.2017.92}}.

\bibitem{Patri2017DiscoveringMW}
Om~Patri, Michael Wojnowicz, and Matt Wolff.
\newblock Discovering malware with time series shapelets.
\newblock In {\em HICSS}, 2017.

\bibitem{DBLP:journals/corr/abs-1708-03346}
Edward Raff and Charles~K. Nicholas.
\newblock Lempel-ziv jaccard distance, an effective alternative to ssdeep and
  sdhash.
\newblock {\em CoRR}, abs/1708.03346, 2017.
\newblock URL: \url{http://arxiv.org/abs/1708.03346}, \href
  {http://arxiv.org/abs/1708.03346} {\path{arXiv:1708.03346}}.

\bibitem{10.1007/978-981-10-6898-0_19}
R.~K. Rahul, T.~Anjali, Vijay~Krishna Menon, and K.~P. Soman.
\newblock Deep learning for network flow analysis and malware classification.
\newblock In Sabu~M. Thampi, Gregorio Mart{\'i}nez~P{\'e}rez, Carlos~Becker
  Westphall, Jiankun Hu, Chun~I. Fan, and F{\'e}lix G{\'o}mez~M{\'a}rmol,
  editors, {\em Security in Computing and Communications}, pages 226--235,
  Singapore, 2017. Springer Singapore.

\bibitem{7413680}
J.~Saxe and K.~Berlin.
\newblock Deep neural network based malware detection using two dimensional
  binary program features.
\newblock In {\em 2015 10th International Conference on Malicious and Unwanted
  Software (MALWARE)}, pages 11--20, Oct 2015.
\newblock \href {http://dx.doi.org/10.1109/MALWARE.2015.7413680}
  {\path{doi:10.1109/MALWARE.2015.7413680}}.

\bibitem{Scofield:2017:FML:3151137.3151142}
Daniel Scofield, Craig Miles, and Stephen Kuhn.
\newblock Fast model learning for the detection of malicious digital documents.
\newblock In {\em Proceedings of the 7th Software Security, Protection, and
  Reverse Engineering / Software Security and Protection Workshop}, SSPREW-7,
  pages 3:1--3:8, New York, NY, USA, 2017. ACM.
\newblock URL: \url{http://doi.acm.org/10.1145/3151137.3151142}, \href
  {http://dx.doi.org/10.1145/3151137.3151142}
  {\path{doi:10.1145/3151137.3151142}}.

\bibitem{shahi2009}
G.S. Shahi, E.F. Pang, and P.P.E. Fong.
\newblock {\em Technology in a Changing World}.
\newblock Lulu Enterprises Incorporated, 2009.
\newblock URL: \url{https://books.google.co.il/books?id=ZTcIAgAAQBAJ}.

\bibitem{webeye2017}
Johann Vierthaler, Roman Kruszelnicki, and Julian Schütte.
\newblock Webeye automated collection of malicious http traffic.
\newblock Technical report, Fraunhofer Research Institute for Applied and
  Integrated Security, November 2017.

\bibitem{DBLP:journals/corr/WangGQ16}
Beilun Wang, Ji~Gao, and Yanjun Qi.
\newblock A theoretical framework for robustness of (deep) classifiers under
  adversarial noise.
\newblock {\em CoRR}, abs/1612.00334v8, Feb 2017.
\newblock URL: \url{http://arxiv.org/abs/1612.00334v8}, \href
  {http://arxiv.org/abs/1612.00334v8} {\path{arXiv:1612.00334v8}}.

\bibitem{7796904}
M.~Wojnowicz, D.~Zhang, G.~Chisholm, X.~Zhao, and M.~Wolff.
\newblock Projecting ``better than randomly'': How to reduce the dimensionality
  of very large datasets in a way that outperforms random projections.
\newblock In {\em 2016 IEEE International Conference on Data Science and
  Advanced Analytics (DSAA)}, pages 184--193, Oct 2016.
\newblock \href {http://dx.doi.org/10.1109/DSAA.2016.26}
  {\path{doi:10.1109/DSAA.2016.26}}.

\bibitem{ndss18}
Weilin Xu, David Evans, and Yanjun Qi.
\newblock Feature squeezing: Detecting adversarial examples in deep neural
  networks.
\newblock In {\em NDSS}, 2018.

\bibitem{10.1007/978-3-319-73830-7_6}
Zhiwu Xu, Cheng Wen, Shengchao Qin, and Zhong Ming.
\newblock Effective malware detection based on behaviour and data features.
\newblock In Meikang Qiu, editor, {\em Smart Computing and Communication},
  pages 53--66, Cham, 2018. Springer International Publishing.

\bibitem{Yan2018scn}
Jinpei Yan, Yong Qi, and Qifan Rao.
\newblock Detecting malware with an ensemble method based on deep neural
  network.
\newblock {\em Security and Communication Networks}, Feb 2018.

\bibitem{7966342}
M.~Yousefi-Azar, V.~Varadharajan, L.~Hamey, and U.~Tupakula.
\newblock Autoencoder-based feature learning for cyber security applications.
\newblock In {\em 2017 International Joint Conference on Neural Networks
  (IJCNN)}, pages 3854--3861, May 2017.
\newblock \href {http://dx.doi.org/10.1109/IJCNN.2017.7966342}
  {\path{doi:10.1109/IJCNN.2017.7966342}}.

\bibitem{DBLP:journals/corr/abs-1708-08042}
Songqing Yue.
\newblock Imbalanced malware images classification: a {CNN} based approach.
\newblock {\em CoRR}, abs/1708.08042, 2017.
\newblock URL: \url{http://arxiv.org/abs/1708.08042}, \href
  {http://arxiv.org/abs/1708.08042} {\path{arXiv:1708.08042}}.

\bibitem{Yuxin2017}
Ding Yuxin and Zhu Siyi.
\newblock Malware detection based on deep learning algorithm.
\newblock {\em Neural Computing and Applications}, Jul 2017.
\newblock URL: \url{https://doi.org/10.1007/s00521-017-3077-6}, \href
  {http://dx.doi.org/10.1007/s00521-017-3077-6}
  {\path{doi:10.1007/s00521-017-3077-6}}.

\bibitem{7847046}
Y.~Zhang, Q.~Huang, X.~Ma, Z.~Yang, and J.~Jiang.
\newblock Using multi-features and ensemble learning method for imbalanced
  malware classification.
\newblock In {\em 2016 IEEE Trustcom/BigDataSE/ISPA}, pages 965--973, Aug 2016.
\newblock \href {http://dx.doi.org/10.1109/TrustCom.2016.0163}
  {\path{doi:10.1109/TrustCom.2016.0163}}.

\bibitem{8029425}
Y.~Zhang, C.~Rong, Q.~Huang, Y.~Wu, Z.~Yang, and J.~Jiang.
\newblock Based on multi-features and clustering ensemble method for automatic
  malware categorization.
\newblock In {\em 2017 IEEE Trustcom/BigDataSE/ICESS}, pages 73--82, Aug 2017.
\newblock \href {http://dx.doi.org/10.1109/Trustcom/BigDataSE/ICESS.2017.222}
  {\path{doi:10.1109/Trustcom/BigDataSE/ICESS.2017.222}}.

\bibitem{Oscar2017}
Morten~Oscar Østbye.
\newblock Multinomial malware classification based on call graphs, May 2017.

\end{thebibliography}

\documentclass[a4paper]{article}

\usepackage[english]{babel}
\usepackage[utf8x]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{multirow}
\usepackage{pifont}

\usepackage[a4paper,top=3cm,bottom=2cm,left=3cm,right=3cm,marginparwidth=1.75cm]{geometry}

\usepackage{amsmath}
\usepackage{graphicx}
\usepackage[colorinlistoftodos]{todonotes}
\usepackage[colorlinks=true, allcolors=blue]{hyperref}
\usepackage{authblk}
\usepackage{url}

\title{Microsoft Malware Classification Challenge}
\author[1]{{\normalsize Royi Ronen}}
\author[2]{{\normalsize Marian Radu\footnote{This work was done while this author was with Microsoft.}}}
\author[1]{{\normalsize Corina Feuerstein}}
\author[3]{{\normalsize Elad Yom-Tov}}
\author[4]{{\normalsize Mansour Ahmadi}}
\affil[1]{{\normalsize Microsoft}}
\affil[2]{{\normalsize CrowdStrike}}
\affil[3]{{\normalsize Microsoft Research}}
\affil[4]{{\normalsize Northeastern University}}

\begin{document}
\date{{\small \{royir,corinaf,eladyt\}@microsoft.com, marian.radu@crowdstrike.com, m.ahmadi@northeastern.edu}}
\maketitle

\begin{abstract}
The Microsoft Malware Classification Challenge was announced in 2015 along with a publication of a huge dataset of nearly 0.5 terabytes, consisting of disassembly and bytecode of more than 20K malware samples. Apart from serving in the Kaggle competition, the dataset has become a standard benchmark for research on modeling malware behaviour. To date, the dataset has been cited in more than 50 research papers. 
Here we provide a high-level comparison of the publications citing the dataset. The comparison simplifies finding potential research directions in this field and future performance evaluation of the dataset.
\end{abstract}

\section{Introduction}

In recent years, the malware industry has become a large and well-organized market \cite{shahi2009}. Well funded, multi-player syndicates heavily invest in technologies and capabilities built to evade traditional protection, requiring anti-malware vendors to develop counter-mechanisms for finding and deactivating them. In the meantime, they inflict significant financial loss to users of computer systems.

One of the major challenges that anti-malware software faces today are the vast amounts of data which needs to be evaluated for potential malicious intent. For example, Microsoft's real-time anti-malware detection products executes on over 600M computers worldwide \cite{msrt18}. This generates tens of millions of daily data points to be analyzed as potential malware. One of the main reasons for these high volumes of different files is that in order to evade detection, malware authors introduce polymorphism to the malicious components. This means that malicious files belonging to the same malware ``family'', with the same forms of malicious behavior, are constantly modified and/or obfuscated using various tactics, so that they appear to be many different files.

A first step in effectively analyzing and classifying such a large number of files is to  group them and identify their respective families. In addition, such grouping criteria may be applied to new files encountered on computers in order to detect them as malicious and of a certain family. To facilitate research in this area, especially in the development of effective techniques for grouping variants of malware files into their respective families, Microsoft provided the data science and security communities with a malware dataset of unprecedented size. Here we summarize the many uses of this dataset, published to date.

\section{Dataset}

The malware dataset is almost half a terabyte when uncompressed. It consists of a set of known malware files representing a mix of 9 different families. Each malware file has an identifier, a 20 character hash value uniquely identifying the file, and a class label, which is an integer representing one of the 9 family names to which the malware may belong (See Table~\ref{tab:training}). For each file, the raw data contains the hexadecimal representation of the file's binary content, without the header (to ensure sterility).  The dataset also includes a metadata manifest, which is a log containing various metadata information extracted from the binary, such as function calls, strings, etc. This was generated using the IDA disassembler tool. The original question posed to participants was to classify malware to one of the 9 classes. The dataset can be downloaded from the competition website.\footnote{https://www.kaggle.com/c/malware-classification/data} 

\begin{table}[tbp]
\centering
\caption{Malware families in the dataset}
\label{tab:training}
\begin{tabular}{lcc}
\noalign{\smallskip} \hline\noalign{\smallskip} 
Family Name    & \# Train Samples & Type \\
\hline\noalign{\smallskip} 
Ramnit         & 1541              &     Worm        \\
Lollipop       & 2478              &     Adware        \\
Kelihos\_ver3  & 2942              &     Backdoor        \\
Vundo          & 475               &     Trojan        \\
Simda          & 42                &     Backdoor        \\
Tracur         & 751               &     TrojanDownloader        \\
Kelihos\_ver1  & 398               &     Backdoor        \\
Obfuscator.ACY & 1228              &     Any kind of obfuscated malware       \\
Gatak          & 1013              &     Backdoor       \\
\hline\noalign{\smallskip} 
\end{tabular}
\end{table}

\section{Citations Comparison}

Since the end of the competition in April 2015, more than 50 research papers and thesis works cited the competition and the dataset. 
Among the citations, several papers are not in English, which we are unable to read \cite{russian2016,korean2016,french2015,NODE07303202}. 
The remaining articles can be divided into two principal classes. The first category of papers referenced the challenge to either perform an abstract comparison or highlight the importance of machine learning for malware classification in industry, where the size of data is huge \cite{7413680,DBLP:journals/corr/GarciaM16,imran2016,DBLP:journals/corr/WangGQ16,DBLP:journals/corr/GaoWQ17,8062202,ndss18,Scofield:2017:FML:3151137.3151142,10.1007/978-981-10-7080-8_17,DBLP:journals/corr/abs-1708-08042,webeye2017,Hansen2017,10.1007/978-3-319-60876-1_10,Oscar2017,Fields2016,Fowler2015,8102915,10.1007/978-3-319-73830-7_6}. 
Papers in the second category performed partial or complete evaluation on the dataset to verify the effectiveness and/or efficiency of their proposed approach for various tasks. We list the papers of the second category in Table~\ref{tab:comparison} sorted by the publication date. Moreover, we summarize the main contribution or focus of each paper to make higher level clusters. Feature engineering, feature selection/fusion, being scalable, being robust, malware authorship attribution, detecting concept drift, performing a measurement, similarity hashing, classification techniques and deep learning are the major contributions of the papers. The diversity of the contributions has made the dataset a benchmark for various tasks, helping researchers provide a standard for evaluation and comparison. 

\begin{table}[tbp]
\centering

\begin{tabular}{ccl|c}
\hline\noalign{\smallskip} 
Month & Year & Method & Focus/Contribution of Research \\
\noalign{\smallskip}\hline\noalign{\smallskip}
Mar & 2016 & Ahmadi et al. \cite{Ahmadi:2016:NFE:2857705.2857713} & Feature Engineering,Feature Fusion,Being Scalable \\
May & 2016 & Drew et al. \cite{7527757} & Feature Engineering,Being Scalable \\
Jul & 2016 & Hu et al. \cite{7523365} & Being Scalable \\
Jul & 2016 & Narayanan et al. \cite{7856826} & Feature Engineering \\
Jul & 2016 & Celik et al. \cite{DBLP:journals/corr/CelikMIPS16} & Being Robust  \\
Aug & 2016 & Zhang et al. \cite{7847046} & Being Scalable,Classification Techniques \\
Sep & 2016 & Bhattacharya et al. \cite{DBLP:journals/corr/BhattacharyaMBC16} &  Being Scalable,Classification Techniques \\
Oct & 2016 & Dinh et al. \cite{7723750} & Classification Techniques \\
Oct & 2016 & Wojnowicz et al. \cite{7796904} &  Feature Reduction \\
Nov & 2016 & Borbely \cite{Borbely2016} & Clustering Techniques \\
Dec & 2016 & Burnaev et al. \cite{7836677} & Classification Techniques  \\
Dec & 2016 & Alrabaee et al. \cite{10.1007/978-3-319-51966-1_17} & Malware Authorship Attribution \\
\noalign{\smallskip}\hline\noalign{\smallskip}
Jan & 2017 & Drew et al. \cite{Drew2017} & Being Scalable,Classification Techniques  \\
Jan & 2017 & Patri et al. \cite{Patri2017DiscoveringMW} & Classification Techniques  \\
Mar & 2017 & Hassen et al. \cite{Hassen:2017:SFC:3029806.3029824} & Feature Engineering,Being Scalable \\
Mar & 2017 & Celik et al. \cite{Celik:2017:FCP:3041008.3041018} & Being Robust  \\
May & 2017 & Yousefi-Azar et al. \cite{7966342} & Feature Engineering \\
Jun & 2017 & Kebede et al. \cite{8268747} & Deep Learning \\
Jul & 2017 & Yuxin et al. \cite{Yuxin2017} & Deep Learning \\
Aug & 2017 & Zhang et al. \cite{8029425} & Clustering Techniques \\
Aug & 2017 & Jordaney at al. \cite{203684} & Detecting Concept Drift  \\
Aug & 2017 & Raff et al. \cite{DBLP:journals/corr/abs-1708-03346} & Similarity Hashing \\
Oct & 2017 & Kim et al. \cite{Kim2017} & Deep Learning \\
Nov & 2017 & Rahul et al. \cite{10.1007/978-981-10-6898-0_19} & Deep Learning \\
Dec & 2017 & Bagga \cite{Bagga2017} & Measurement and Comparison  \\
Dec & 2017 & Gsponer et al. \cite{10.1007/978-3-319-71246-8_3} & Classification Techniques \\
Dec & 2017 & Hassen et al. \cite{Hassen2017} & Feature Engineering \\
Dec & 2017 & Fan et al. \cite{8064195} & Being Scalable \\
Dec & 2017 & Kim \cite{10.1007/978-981-10-7605-3_215} & Deep Learning  \\
\noalign{\smallskip}\hline\noalign{\smallskip}
Jan & 2018 & Hwang et al. \cite{fselection2018} & Feature Selection \\
Feb & 2018 & Yan et al. \cite{Yan2018scn} & Deep Learning \\
Feb & 2018 & Kreuk et al. \cite{2018arXiv180204528K} & Adverserial Examples,Deep Learning \\
Feb & 2018 & Hassen et al. \cite{2018arXiv180204365H}  & Open Set Recognition \\
\noalign{\smallskip}\hline
\noalign{\smallskip}
\end{tabular}

\smallskip
\caption{A comparison between research papers that have performed partial or complete evaluation on Microsoft malware classification challenge dataset.}
\label{tab:comparison}
\end{table}

\section{Conclusion and Future Directions}

In this paper, we provide a short description of the characteristics of the Microsoft Malware Classification Challenge dataset. This dataset is becoming a standard dataset with more than 50 papers citing it. We enumerated these references as much as possible and compared their main contributions with respect to the dataset. The comparison helps the understanding of what the existing contributions are, and what  the potential research directions can be.

The authors aim to keep the reference table updated. We encourage the community to cite this paper when using the dataset, and update us about such work so it can be added to this paper.

\bibliographystyle{plainurl}
\bibliography{sample}

\end{document}